------------------------------------------------------------------------------ Time Accounting: (June 5 -- June 12) ------------------------------------------------------------------------------ Hall Availability: 96% Scheduled Program: 161 hours Sched. hours minus accel config change: 134 hours ABU: 80 hours (60%) (ABUs computed from Sat--Monday are closer to 40%) ------------------------------------------------------------------------------ Overview ------------------------------------------------------------------------------ The Hall made the switch from the He-elastic experiment (E04-018) to GammaPP (E03-101) on Friday, June 8. This was one day later than the formal schedule to allow E04-018 to complete their He4 running. Changeover activities are listed in the next section. GammaPP's focus in on measuring the hard photodisintegration of the pp pair inside the He3 nucleus. A typical kinematic point involves setting both HRSs to equal angles representing +-90 degrees in the center of mass frame to detect both ejected protons in coincidence (the neutron is a spectator). The resulting He3(g,pp)n cross sections allow a strong test of several quark models over the photon energy range 1--5 GeV. Secondary physics goals include measuring recoil polarization on carbon and copper. As of June 12 the experiment has been progressing quite well. Hall availability has been very high. The experiment has collected data at beam energies of 1 and 2 GeV and begun taking data at 3 GeV. Polarization transfer data were taken at 1 GeV but MCC had to back out the Wien angle changes at 2 GeV due to unexpected and severe beam transport problems for Hall C that originated in the injector region. Those problems have (understandably) made the other Halls very nervous about Hall A's tentative request to complete the polarization measurement at 1.6 GeV (2 pass @ ~800 MeV/pass) starting June 19. The PD (Riad) has requested that the polarization request (ie. Wien angle change) be formally cleared with Dennis Skopik as soon as possible. The Hall A Moeller quads were repaired (coolant hoses replaced) and was recommissioned on June 9. Moeller results for the 1 pass beam (1 GeV): target 5: -42.3+/-0.2% target 3: -41.2+/-0.3% This is consistent with expected polarization off the bulk cathode. MCC continues to struggle with machine stability and procedural issues that prolong pass changes, energy changes, etc beyond the time allotted. There have also been communication issues with MCC that have aggravated all three Halls (ie. beam studies and impromptu test plans being implemented or scheduled with little or no notice to the Halls). Note that the communication has been quite good over the last 3-4 days so hopefully this problem has been resolved. ============================================================================== ------------------------------------------------------------------------------ Changeover activity on Friday, June 8 ------------------------------------------------------------------------------ - Moeller quad coolant hoses replaced (all were very brittle) - Moeller system is now functional - Target changed from He4 -> He3 - S0 installed in RHRS - FPP recommissioned in LHRS (only rear chambers are installed) - Both arms configured for positive polarity - Radiator recommissioned - 'manual' motion control of the stepper motor hooked up - telephoto camera set up to view radiator position - radiator temperature readback via software (EPICs) fixed - Arne replaced dead compact-flash card in Calorimeter IOC that handles the temp. readout ============================================================================== ------------------------------------------------------------------------------ Problems pending (as of June 12) ------------------------------------------------------------------------------ - No physics-critical issues to report - 60Hz noise on linear fan-in unit used with Moeller DAQ - sub-optimal, but system still functions - Eugene is looking for replacement - LHRS detector hut door control issues - one control module won't store settings between power-cycles anymore - 'good' controller swapped to left door - current work-around: left door works, right door disabled - 1HA03B harp scanner not working with current beam position - The beam must be steered to -3.5mm in horizontal, +1mm vertical (BPM coordinates) to be centered on the target ladder. One of the Harp wires doesn't cross the beam at those coords (apparently). - Workaround: use 1HA03A - Still need to request MCC to do the scan as the Hall A software crashes - It would be good to either a) have the Hall A Harps under the formal control of MCC, or b) if Hall A wants to 'drive' then we should trash the current Hall A software and request (limited?) access to the much more functional MCC Harp control software. - Fix zoom on beam-right Hall camera (techs aware, need parts) ============================================================================== ------------------------------------------------------------------------------ Daily Summary ------------------------------------------------------------------------------ ------------------------------------------------------------------------------ # Tuesday, June 5 --- Running 68 fm^-2, E_beam=4.137 GeV, He4 ------------------------------------------------------------------------------ ------------------------------------------------------------------------------ # Wednesday, June 6 --- Running 68 fm^-2, E_beam=4.137 GeV, He4 ------------------------------------------------------------------------------ ------------------------------------------------------------------------------ # Thursday, June 7 --- Running 68 fm^-2, E_beam=4.137 GeV, He4 ------------------------------------------------------------------------------ ------------------------------------------------------------------------------ # Friday, June 8 --- Changeover to E03-101 ------------------------------------------------------------------------------ - changeover to E03-101 - target change: He4 -> He3 - scheduled start (schedule: 0900 -- 1800; real: 0900 -- - restore beam (schedule: 1800 -- 2200 - required pass change: 4-> 1 pass - "Moeller quad optics" restored - polarized beam to Hall A ------------------------------------------------------------------------------ # Saturday, June 9 --- Running E_beam = 1.077 GeV ------------------------------------------------------------------------------ - Got first CW beam around 0200 and began commissioning - Started production running around 0500 - iocha14 (LHRS magnet controls) died - Jack needed to swap out the VME crate (PS problem) - Jack swapped control modules on LHRS shield hut doors - left door now working reliably, right door disabled - Eugene noted the Wien angle had been computed incorrectly for 1 pass to our Hall (-50 deg) and informed MCC - MCC worked out the correct angle (36.7 degrees), changed angle, and retuned optics - Moeller Results: target 5: -42.3+/-0.2% target 3: -41.2+/-0.3% - ROC2 locked up, reboot required controlled access - Noticed that the ozone monitor had been deadlined since 3pm Saturday - fix deferred until Monday ------------------------------------------------------------------------------ # Sunday, June 10 --- Running E_beam = 1.077, 2.097 GeV ------------------------------------------------------------------------------ - Pass changes (scheduled: 0830--1230; real: 0830--1900) Hall A 1 -> 2 Hall A 2 -> 3 - Switched to 2GeV kinematics - Reasonable production running - fairly high trip rate (~15--20/hour) through to Monday morning ------------------------------------------------------------------------------ # Monday, June 11 --- Running E_beam = 2.097, 3.177 GeV ------------------------------------------------------------------------------ - Techs (Jack, Heidi, Mark) fixed the ozone readback - IOC (iocha12) was not dead, but the software task that readout the ozone monitor was not running. There was an undocumented(?) change by the "software people" that caused the problem. One the change was backed out we got our readback back. I don't know why the problem chose to show up at 3pm on Saturday... - MCC changed the Wien angle back to 0 degrees and did a few hours of injector work in an attempt to stabilize beam to Hall C - Pass change Hall A 2 -> 3 (only took ~90 minutes, great!) - Prompted by Hall C request to shift their pass change to Wednesday (previous plan had us both changing pass on Tuesday). We had finished what we had to do at 2 GeV (so we elected to go to 3 pass early. ------------------------------------------------------------------------------ # Tuesday, June 12 --- Running E_beam = 3.177 GeV ------------------------------------------------------------------------------ - problem with nitrogen level monitor on RHRS (Q1) - Heidi and Mark taking a look at it during opportunistic access this morning -- Fixed - ozone readback still seems flaky -- there is now bit-flicker, but it's still a flatline with or w/o beam... ------------------------------------------------------------------------------ # Wednesday, June 13 --- Running E_beam = 3.177 GeV ------------------------------------------------------------------------------ - Hall C pass change from 0900--1100 - Folts, Heidi, EHS personnel took opportunistic access to get more info regarding the unexpectedly flat ozone readback from the Hall A meter - Beam studies from 1100--1900 - "High current optimization" to Hall A and C - we will have Carbon target in place, should expect up to 100 uA - Took some 'parasitic' radiator out, C target data during beam studies - high trip rate, unstable beam, might not be useful